a first faces the demand curve, Q=180-5p, and has the cost equation, c=100+24Q. what is the equation for the firm's marginal revenue?
Answer #1

Answer

Inverting demand curve to inverse demand curve

Q=180-5P

5P=180-Q

P=36-0.2Q

total revenue =P*Q

=(36-0.2Q)*Q

TR=36Q-0.2Q^2

marginal revenue is the change in total revenue. a change in function found by differentiation.

MR=dTR/dQ=36-0.4Q ....... the first differential of TR function concerning Q.

MR=36-0.4Q

the marginal revenue function is

MR=36-0.4Q